# Adanonymizer: Interactively Navigating and Balancing the Duality of Privacy and Output Performance in Human-LLM Interaction

### v1

@@ Abstract.
Current Large Language Models (LLMs) cannot support users to precisely balance privacy protection and output performance during individual consultations. We introduce Adanonymizer, an anonymization plug-in that allows users to control this balance by navigating a trade-off curve. A survey (N=221) revealed a privacy paradox, where users frequently disclosed sensitive information despite acknowledging privacy risks. The study further demonstrated that privacy risks were not significantly correlated with model output performance, highlighting the potential to navigate this trade-off. Adanonymizer normalizes privacy and utility ratings by type and automates the pseudonymization of sensitive terms based on user preferences, significantly reducing user effort. Its 2D color palette interface visualizes the privacy-utility trade-off, allowing users to adjust the balance by manipulating a point. An evaluation (N=36) compared Adanonymizer with ablation methods and differential privacy techniques, where Adanonymizer significantly reduced modification time, achieved better perceived model performance and overall user preference.
@@ 1. Introduction
The potential information disclosure during users’ interaction was more severe compared with traditional text input, as evidenced in the past literature (Zhang et al., 2024a). However, in contrast to to the rapid adoption and increasing risk of LLMs, the privacy regulation of its input and output lacked systematic investigation. Traditional anonymization techniques like paraphrasing (Staufer et al., 2024) lacked the consideration of model performance. Differential privacy techniques (Fernandes et al., 2019) lack the transparency and users’ control. Besides, the text information sensitivity of differential privacy techniques was decided by users’ tasks and were often opaque (Franzen et al., 2022). Encryption based techniques (Lin et al., 2024a) faced a high computational cost and also an accuracy loss. In fact, the privacy protection and model performance was long argued as a “trade-off” relation (Frikha et al., 2024; Staab et al., 2024), where high privacy protection requirements may resulted in unusable model performance. The constraint between privacy protection and model performance significantly hindered the adoption of text-based privacy protection techniques (Lison et al., 2021).
@@
Collaborative interaction allows users and LLMs to jointly modify content, enabling users to better control privacy protection and model output performance according to their personalized preferences (Asthana et al., 2024; Zhang et al., 2024a). In response to these challenges, this work addresses the core research question: ‘How can privacy protection and model output performance be balanced through collaborative interaction between humans and AI in human-LLM Interactions?’ Specifically, we focus on, (1) explore users’ perceptions of privacy protection and model output performance, and (2) design a collaborative interaction system that balances both.
@@
We initiated the first study (N=221) to examine users’ privacy perceptions regarding their input during human-LLM interactions. Through designing Personal Information (PI) disclosure vignettes under the specific scenario as the questionnaire, we measured participants’ perceived privacy risk, utility and occurrence frequency for different PI when they were included in the input of human-LLM interaction. The privacy risks were assessed with respect to two emerging and correlated privacy risks of private input data: information leakage risk and retraining risk. The results showed that users still suffered from “Privacy Paradox” (Nissenbaum, 2011), who input specific types of private data frequently despite recognizing the privacy. Users were found to care about their convenience usually without noticing the potential privacy risk. We also found no significant correlation between privacy risks and perceived model output performance degradation when deleting private information. Personal health and identification data presented higher privacy risks and a significant impact on model performance, while basic personal information exhibited lower privacy risks and minimal effect on model output. This discrepancy between privacy and utility informed the design of Adanonymizer.
@@
We introduce Adanonymizer, a collaborative anonymization technique with a color palette-like design, allowing users to balance privacy and output performance (see Figure 1). Following the “notice and choice” principle (Nissenbaum, 2011), Adanonymizer visualizes the privacy-utility trade-off as a curve on a two-dimensional plane, allowing users to intuitively explore the trade-off by selecting points below or near the curve and receiving immediate feedback, similar to adjusting a color palette. Users can click near the curve to find candidates that balance privacy and model output performance more effectively. Adanonymizer anonymizes different types of information based on users’ privacy and performance preferences, as well as the results from Study 1, with the option for further modifications if necessary. Adanonymizer identifies personal information through a prompt-based approach and selectively anonymizes it according to the chosen proportion.
@@
A usability evaluation study with 36 users compared Adanonymizer with its ablation techniques and a Differential Privacy (DP)-based baseline technique. We selected the personal consultation task under three different context (Zhang et al., 2024a) (work-related, academic-related and life-related) where users’ input may frequently contain private personal information. Adanonymizer outperformed other baseline techniques in terms of modification time, users’ perceived model output performance and perceived satisfaction. Besides, users also favored Adanonymizer for the better experience, the higher perceived control and transparency.
@@
  • We evaluated users’ privacy risk of PI when interacting with LLMs through a questionnaire study (N=221). Participants have privacy paradox when including PI in their input. Additionally, the privacy risk and the influence on model’s performance of each PI varied, paving the foundation for balancing privacy and model performance.
  • We proposed Adanonymizer, which mapped the abstract privacy protection and model performance control onto a 2D color palette metaphor. Adanonymizer supported users to select their preferred balance point through clicking on the palette and immediately getting the feedback.
  • An evaluation study (N=36) comparing Adanonymizer with two ablation techniques and another differential privacy based baseline showed Adanonymizer outperformed other techniques in terms of modification time, perceived output performance and comparable privacy protection capability. Adanonymizer was also praised for its transparency and control-ability.
@@ 2.1. Privacy Risk in LLM-based Conversational Assistants
We detail two distinctive privacy challenges in LLM-based conversational assistants (CAs): 1) memorization risks and 2) the human-like interactions that nudge users to disclose more information.
LLM-based CAs are built primarily on LLMs. To enhance their conversational capabilities, LLMs require vast amounts of data for their training, often encompassing user interaction data (Pahune and Chandrasekharan, 2023). However, a side effect of this data-driven nature of LLMs is the unintentional memorization of the training data, which may include user inputs containing personally identifiable information (PII) (Brown et al., 2022; Peris et al., 2023). This PII can sometimes be unintentionally reproduced in the generated output. For example, despite safety precautions, ChatGPT can inadvertently disclose PII through carefully crafted prompts.
Users engage with LLM-based CAs through natural language, which is traditionally reserved for human-to-human communication. This can lead them to perceive these agents as human-like. Studies suggested that anthropomorphizing can increase users’ trust in CAs and more user information disclosure (Ischen et al., 2020; Kim and Sundar, 2012). Anthropomorphizing can inflate users’ perceptions of the CA’s competencies, fostering undue confidence, trust, or expectations in these agents (Kim and Sundar, 2012; Złotowski et al., 2015). As trust grows, users might be more inclined to share private information, even in sensitive contexts (Kim and Sundar, 2012; Waldman, 2018; Złotowski et al., 2015). Anthropomorphization may amplify the risks of users yielding effective control by trusting CAs unquestioningly. Moreover, more private information may be revealed when CAs leverage psychological effects, such as nudging or framing (Weidinger et al., 2021). In comparison, this study provides the first empirical evidence supporting the notion that LLMs’ human-like text generation can prompt more sensitive disclosure behaviors in certain contexts.
Generalized to AI technologies, previous literature (Lee et al., 2024c) classified the risks into: invasion risks (Solove, 2005; Otrel-Cass et al., 2010; Ping et al., 2018; Milmo, 2021), data collection risks (Solove, 2005), data dissemination risks (Ayyub, 2018; Levin, 2017; Burgess, 2021) and data processing risks (Solove, 2005; Simonite, 2018; Pannett, 2022; Pearl, 2010; Stark and Hutson, 2021). Intrusion risks includes actions that disturb one’s solitude in physical space (Solove, 2005), such as personalized advertisements. In addition, surveillance is often supported by AI technologies (Otrel-Cass et al., 2010; Ping et al., 2018), and ubiquitous sensors (Milmo, 2021). Data collection risks “create disruption based on the process of data gathering” (Solove, 2005), which exacerbated surveillance risks (Solove, 2005). Data processing risks result from the use, storage, and manipulation of personal data (Solove, 2005; Simonite, 2018). Aggregation, which is a form of data processing risk refer to combining various pieces of data to make inferences beyond what is explicitly captured in those data (Solove, 2005). Relatedly, debunked pesudosciences like phrenology and physiognomy falsely claim to infer personality or character from outer appearance (Stark and Hutson, 2021). Data collected may also be subjected to secondary use without consent (Solove, 2005; Pannett, 2022). Data dissemination risk contained exposure, distortion, disclosure, increased accessibility, and others. Via generative techniques, AI can create, reconstruct, manipulate content (Levin, 2017; Burgess, 2021) or cause harassment (Ayyub, 2018). These risks may be exacerbated upon the privacy leakage of users’ input.
@@ 2.3. Privacy-utility Trade-off in Privacy Regulation of Text
In this section, we applied Contextual Integrity theory (Nissenbaum, 2004) to survey the sensitivity and perceived frequency of different types of potentially disclosed information.
@@ 3.1. Recruitment and Participants
We recruited 221 Chinese participants (123 females, 98 males) through distributing the questionnaire on an online recruiting platform  [^1] https://wjx.cn, last accessed at Sep 12nd, 2024. To ensure diversity, we distributed the questionnaire in different time across three days. Participants were required to be between 18 and 65 years old. 141 participants were 26 to 35 years old with 47 participants 36 to 45 years old. 56 were with a manufacturing occupation including mechanical engineers, factory works, etc. 45 were from information technology occupation such as software development engineers and system administrators. Others are from human resources, retail, service industry or other fields. 171 participants held a bachelor degree, and 32 had a master’s degree or higher. The detailed demographics of participants were shown in Table 2 in the appendix. Each participant who completed the experiment received 20 RMB as compensation according to the local wage standard.
@@ 3.2. Scope of Personal Information
Personal data, as defined in General Data Protection Regulation (GDPR), refers to “any information related to an identified or identifiable natural person”. In designing scenarios and personal information (PI) types, we followed Contextual Integrity (Nissenbaum, 2004), China’s classification criteria [^2] https://www.tc260.org.cn/upload/2021-12-31/1640948142376022576.pdf and the previous studies (Richthammer et al., 2014, 2013; Chua et al., 2021; Malgieri, 2016) in designing PI categories and types. We compared and synthesized classifications from the aforementioned literature. Due to varying interpretations of personal and privacy information, the information categories identified in different studies are inconsistent. These studies often focus on specific contexts, such as social media applications. To address this, we aggregated information categories from multiple sources and organized them into 14 broader categories, resulting in a comprehensive classification.
We selected 14 categories of information: personal basic information, personal identification information, online identity information, personal health and physiological information, personal medical information, personal education and employment information (including personal education information and personal employment information), personal property information (including financial account information, personal transaction information, personal asset information, personal loan information), identity verification information, personal communication information, contact information, personal internet records, personal location information, personal exercise information, and other personal information. The detailed information types were determined and described in the appendix Section E.
@@ 3.3. Experiment Design
We conducted a one-factor within-subjects study with PI as the sole factor. We examined 14 categories of PI, including personal identity and exercise information, along with 74 specific types, such as name and age, as outlined in the previous subsection.
To ensure participants understood the privacy risk, we first provided descriptions of privacy risks and each PI in the questionnaire. The two types of privacy risk described were information leakage and model retraining risks. Information Leakage Risk refers to the potential disclosure of sensitive information during usage, such as business data or academic assignments, to malicious third-parties. Retraining Risk arises when users’ sensitive input data is used by service providers to improve models through further training. We presented several examples classified by scenario to illustrate each PI, with each example including multiple PI types. We additionally added test questions before the rating entries to ensure the understanding of participants. We measured privacy risk, perceived usage frequency, and perceived utility degradation using a 7-point Likert scale. In total, there were 222 Likert scale entries (3 factors $\times$ 74 PI types). We manually crafted a few cases for illustration of the perceived utility degradation. Detailed information about the questionnaire can be found in Section B.
@@ 3.4. Pilot Studies
Due to the complexity of the interaction with LLMs, traditional classification and summarization of PI types could hardly cover all the PI types when interacting with LLMs. Thus we first conducted three rounds of pilot studies, each involving 10 participants to derive at a comprehensive set of PI types specific to LLM interactions. We added additional subjective open-ended questions to allow participants to suggest additional sensitive information they might disclose to LLMs. To encourage full disclosure and compensate for their effort, participants were awarded an additional 1RMB for each type of sensitive information disclosed. We followed data saturation theory (Fusch Ph D and Ness, 2015), ceasing where all participants could not think out new information that they could potentially disclose.
In the pilot study, we found no additional effective information that users would disclose. This drove our decision to include no additional items. We also revised the explanations of privacy risks and data descriptions based on feedback from pilot participants. In the final round of pilot study, participants demonstrated a clear understanding of the potential harms and privacy risks, as well as the scenarios provided, indicating no further misunderstandings.
@@ 3.5. Procedure
Before the experiment, participants were briefed on its purpose and the potential risks involved. They were then asked to sign a consent form before completing the questionnaire. They could quit the experiment at any time. The questionnaire contained 222 questions in total on average, which cost participants 1756 seconds on average (SD=710). No one dropped the experiment. To ensure the questionnaire’s quality, we inserted 4 questions that checked whether users understand private information in the questionnaire. 9/221 (4.07%) participants were filtered out, leaving 212 valid participants in the end.
@@ 3.6. Results
This section presents the perceived privacy risk ratings and perceived frequency distributions for different information types, followed by an analysis of their relationship. Additionally, we examine the relationship between perceived privacy risk ratings and the perceived utility of these information types.
@@ 3.6.1. Perceived Privacy Risk and Usage Frequency of Different Scenarios and PI
Table 1 and Figure 2 presented the privacy risks and appearance frequencies with different scenarios and PI types separately. Our findings indicate that participants were most concerned about Identity Verification Information (M=5.86), while Personal Exercise Information (M=3.39) attracted the least attention. The top three perceived high-risk categories—Identity Verification Information (M=5.86), Online Identity Identifier Information (M=5.74), and Personal Property Information (M=5.57)—likely rank highest because their exposure could lead to significant financial or personal loss. Conversely, Personal Exercise Information (M=3.39), Personal Basic Information (M=4.32), and Personal Health Status and Physiological Information (M=4.94) were rated as lower risk, possibly due to their frequent use and the minimal harm caused by their disclosure. In terms of appearance frequency, Personal Basic Information (M=4.82) appeared most often, followed by Online Identity Identifier Information (M=4.56) and Personal Education and Work Information (M=4.63). The differences in perceived risk and frequency may potentially allow users to select which information to anonymize based on their privacy risk threshold and usage patterns.

### v2

@@ Abstract.
Current Large Language Models (LLMs) cannot support users to precisely balance privacy protection and output performance during individual consultations. We introduce Adanonymizer, an anonymization plug-in that allows users to control this balance by navigating a trade-off curve. A survey (N=221) measured users’ perceived risks and importance of the information for the tasks on nuanced information types. The study also demonstrated that privacy risks were not significantly correlated with model output performance, highlighting the potential to navigate this trade-off. Adanonymizer normalizes privacy and utility ratings by type and automates the pseudonymization of sensitive terms based on user preferences, significantly reducing user effort. Its 2D color palette interface visualizes the privacy-utility trade-off, allowing users to adjust the balance by manipulating a point. The privacy-utility mapping and visualization were set through meticulous iterations and testings. An evaluation (N=36) compared Adanonymizer with ablation methods and differential privacy techniques, where Adanonymizer significantly reduced modification time, achieved better perceived model performance and overall user preference.
@@ 1. Introduction
The potential for information disclosure during user interactions is more pronounced with LLMs compared to traditional text input methods, as highlighted in previous studies (Zhang et al., 2024a). However, despite the rapid adoption of LLMs and the associated rise in privacy risks, systematic investigations into privacy regulations for user inputs remain insufficient. Traditional anonymization techniques like paraphrasing (Staufer et al., 2024) fail to adequately account for model performance. Differential privacy approaches (Fernandes et al., 2019), while effective, lack transparency and user control. Furthermore, the sensitivity of text information in differential privacy methods is often task-dependent and opaque (Franzen et al., 2022). Encryption-based methods (Lin et al., 2024a) suffer from high computational costs and accuracy trade-offs. The relationship between privacy protection and model performance has long been characterized as a “trade-off” (Frikha et al., 2024; Staab et al., 2024), where stringent privacy requirements often degrade model effectiveness. The trade-off has significantly hindered the widespread adoption of privacy-preserving techniques for text-based models (Lison et al., 2021).
@@
Collaborative interaction allows users and LLMs to jointly modify content, which offers poettnail to solve the above challenge as users are the most important judger of utility and privacy protection. Enabling users to better manage privacy protection and model output performance could better fit for their personalized preferences (Asthana et al., 2024; Zhang et al., 2024a). However, how users could effectively balance privacy protection and model output during text-based human-LLM interaction was less explored. Hence this work addresses the key challenge: How can privacy protection and model output performance be balanced through collaborative interaction between humans and AI in human-LLM Interactions? Specifically, we proposed three research questions:
  • RQ1: How to include users’ privacy-utility trade-off cognition into the system design in a quantitative manner?
  • RQ2: How to design a collaborative interaction interface that effectively and efficiently balances both?
  • RQ3: How would users balance privacy and utility through using this sytem in human-LLM interaction?
@@
To investigate RQ1, we conducted a questionnaire study (N=221) to examine users’ privacy perceptions regarding the input they provide during human-LLM interactions. We designed a set of vignettes on Personal Information (PI) disclosure, framed with specific scenarios, as part of a questionnaire to assess participants’ perceived privacy risks, utility and the frequency with which various types of PI were included in their inputs. Privacy risks were evaluated based on two related concerns: information leakage and retaining risks. The results revealed participants’ “Privacy Paradox” (Nissenbaum, 2011), who input specific types of private data frequently despite recognizing its privacy implications. Users were primarily motivated by convenience, often overlooking the associated privacy risks (Zhang et al., 2024a). We also found no significant correlation between privacy risks and perceived model output performance degradation when private information was excluded. Personal health and identification data presented higher privacy risks and had a notable impact on model performance, while basic personal information posed lower risks and had minimal effect on output quality. This divergence between privacy concerns and utility informed the design of Adanonymizer.
@@
We introduce Adanonymizer, a collaborative anonymization technique with a color palette-like design, allowing users to balance privacy and output performance (see Figure 1). Following the “notice and choice” principle (Nissenbaum, 2011), Adanonymizer visualizes the privacy-utility trade-off as a curve on a two-dimensional plane. The visualization adopts normalized linear mapping for clearer communication and selects continuous panel for smoother control. It allows users to intuitively explore the trade-off by selecting points below or near the curve and receiving immediate feedback, similar to adjusting a color palette. Users can click near the curve to find candidates that balance privacy and model output performance more effectively. Adanonymizer anonymizes different types of information based on users’ privacy and performance preferences, as well as the results from Study 1, with the option for further modifications if necessary. Adanonymizer identifies personal information through a prompt-based approach on local quantized models and selectively anonymizes it according to the chosen proportion.
@@
A usability evaluation study with 36 users compared Adanonymizer with its ablation techniques and a Differential Privacy (DP)-based baseline technique. We selected the personal consultation task under three different contexts (Zhang et al., 2024a) (work-related, academic-related and life-related) where inputs often contain private personal information. Adanonymizer outperformed other baseline techniques in terms of modification time, perceived output performance and perceived satisfaction. Additionally, users favored Adanonymizer for its enhanced experience, perceived control and transparency.
@@
  • We conducted a survey (N=221) on users’ privacy risks in LLM interactions, quantifying the privacy paradox and examining the trade-offs between information sensitivity, usage frequency and importance.
  • We introduced Adanonymizer, employing a novel 2D color palette metaphor to help users balance privacy and performance with immediate feedback, offering new insights into privacy-utility trade-offs.
  • We tested Adanonymizer in real scenarios (N=36), revealing efficient privacy-utility balancing strategies, categorized into three distinct exploration patters, with implications for future privacy-utility design.
@@ 2.1. Privacy Risk in LLM-based Conversational Assistants
LLM-based CAs face privacy risks both during training and inference, with the paper focusing primarily on the latter. During training, these agents rely on LLMs, which are trained using vast amounts of data, including user interactions (Pahune and Chandrasekharan, 2023). While this data-driven approach enhances the agents’ conversational abilities, it introduces LLMs’ memorizing risk which could potentially disclose personally identifiable information (PII) from the training data (Brown et al., 2022; Peris et al., 2023). Despite safety measures, LLMs, such as ChatGPT may inadvertently reveal sensitive information when prompted with carefully constructed querie (Kim et al., 2024).
During inference, as users also shared sensitive data (Zhang et al., 2024a), LLMs may collect these data for target advertising, commercial use or used these data for training. Beyond the LLM-specific training use, other privacy risks within AI literature also threatens user-LLM interaction (Lee et al., 2024c), including invasion, data collection, data dissemination and data processing risks (Solove, 2005; Otrel-Cass et al., 2010; Ping et al., 2018; Milmo, 2021). Invasion risks often involve the disruption of personal solitude, such as through targeted advertising or surveillance enabled by AI (Otrel-Cass et al., 2010; Ping et al., 2018; Milmo, 2021). Data collection risks, which pertain to the process of gathering personal data, can exacerbate surveillance concerns (Solove, 2005). Data processing risks emerge from the use, storage, and manipulation of personal data, including aggregation, where disparate data points are combined to make inferences beyond what is explicitly stated (Solove, 2005). Finally, data dissemination risks, including exposure, distortion and unauthorized use, can lead to privacy violations, particularly when AI-generated content is misused (Levin, 2017; Burgess, 2021; Ayyub, 2018). These risks are further amplified when user input is exposed.
@@ 2.3. Privacy-utility Trade-off in Privacy Regulation of Text
In this section, we apply Contextual Integrity theory (Nissenbaum, 2004) to examine the sensitivity and perceived frequency of various types of potentially disclosed information.
@@ 3.1. Recruitment and Participants
We recruited 221 Chinese participants (123 females, 98 males) through distributing the questionnaire on an online recruiting platform  [^1] https://wjx.cn, last accessed at Sep 12th, 2024. To ensure diversity, we distributed the questionnaire in different time across three days. Participants were required to be between 18 and 65 years old. 141 participants were 26 to 35 years old with 47 participants 36 to 45 years old. 56 were with a manufacturing occupation including mechanical engineers, factory works, etc. 45 were from information technology occupation such as software development engineers and system administrators. Others are from human resources, retail, service industry or other fields. 171 participants held a bachelor degree, and 32 had a master’s degree or higher. The detailed demographics of participants were shown in Table 2 in the appendix. Each participant who completed the experiment received 20 RMB as compensation according to the local wage standard.
@@ 3.2. Scope of Personal Information
Personal data, as defined in General Data Protection Regulation (GDPR), refers to “any information related to an identified or identifiable natural person”. In designing scenarios and personal information (PI) categories, we followed Contextual Integrity (Nissenbaum, 2004), China’s classification criteria [^2] https://www.tc260.org.cn/upload/2021-12-31/1640948142376022576.pdf and prior studies (Richthammer et al., 2014, 2013; Chua et al., 2021; Malgieri, 2016). We synthesized classifications from these sources, noting that varying interpretations of personal and privacy information result in inconsistent categories, often tailored to specific contexts, such as social media applications. To address this, we consolidated information from multiple sources into 14 broader categories, providing a comprehensive classification.
These categories include: personal basic information, identification information, online identity, health and physiological data, medical information, education and employment data (covering both education and employment), financial and property data (including financial accounts, transactions, assets, and loans), records, location data, exercise data, and other personal information. A detailed description of these types is provided in Appendix Section E.
@@ 3.3. Experiment Design
We conducted a one-factor within-subjects study with PI as the sole factor, examining 14 PI categories (e.g., including personal identity, exercise information) and 74 specific types (e.g., name, age) as detailed in the previous subsection.
To ensure participants understood privacy risks, we provided descriptions of information leakage and model retraining risks in the questionnaire.Information Leakage Risk refers to the potential disclosure of sensitive information during usage, such as business data or academic assignments, to malicious third-parties. Retraining Risk arises when users’ sensitive input data is used by service providers to improve models through further training. We presented several examples classified by scenario to illustrate each PI, with each example including multiple PI types. We additionally added test questions before the rating entries to ensure the understanding of participants. We measured privacy risk, perceived usage frequency, and perceived utility degradation using a 7-point Likert scale. In total, there were 222 Likert scale entries (3 factors $\times$ 74 PI types). We manually crafted a few cases for illustration of the perceived utility degradation. Detailed information about the questionnaire can be found in Section B.
@@ 3.4. Pilot Studies
Due to the complexity of LLM interactions, traditional PI classifications do not fully capture all relevant types. To address this, we conducted three pilot studies with 10 participants each, refining the set of PI types specific to LLM interactions. Participants were asked open-ended questions to suggest any additional sensitive information they might disclose, and were compensated with 1RMB for each type disclosed. We followed data saturation theory (Fusch Ph D and Ness, 2015), halting when no new disclosures emerged.
The pilot studies revealed no additional types of information, confirming the comprehensiveness of our set. We also revised privacy risk explanations and data descriptions based on participant feedback. In the final round, participants demonstrated clear understanding of the risks and scenarios, with no further misunderstandings.
@@ 3.5. Procedure
Prior to the experiment, participants were briefed on its purpose and risks, and provided informed consent. They could withdraw at any time. The questionnaire contained 222 items, taking an average of 1756 seconds (SD=710) to complete. No participant dropped out. To ensure quality, we included 4 comprehension checks on PI. As a result, 9 of 221 participants (4.07%) were excluded, leaving 212 valid responses.
@@ 3.6. Results
This section presents the perceived privacy risk ratings and perceived frequency distributions for different information types, followed by an analysis of their relationship, including the link between privacy risk and perceived utility.
@@ 3.6.1. Perceived Privacy Risk and Usage Frequency of Different Scenarios and PI
Table 1 and Figure 2 presented the privacy risks and appearance frequencies with different scenarios and PI types separately. Our findings indicate that participants were most concerned about Identity Verification Information (M=5.86), while Personal Exercise Information (M=3.39) attracted the least attention. The top three perceived high-risk categories—Identity Verification Information (M=5.86), Online Identity Identifier Information (M=5.74), and Personal Property Information (M=5.57)—likely rank highest because their exposure could lead to significant financial or personal loss. Conversely, Personal Exercise Information (M=3.39), Personal Basic Information (M=4.32), and Personal Health Status and Physiological Information (M=4.94) were rated as lower risk, possibly due to their frequent use and the minimal harm caused by their disclosure. In terms of appearance frequency, Personal Basic Information (M=4.82) appeared most often, followed by Online Identity Identifier Information (M=4.56) and Personal Education and Work Information (M=4.63). These differences suggest that users may choose to anonymize information based on their privacy risk tolerance and usage patterns.